# references.bib

### v3

@@ Abstract
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
II Related Works
SDE in Medical Endoscopy. SDE in endoscopic images aids in precise navigation, accurate lesion assessment, and improved surgical planning. Given the impracticality of using depth cameras or LiDAR in such settings, most datasets depend on correlating endoscopic views with pre-operative CT scans or generating sparse point clouds via calibrated stereo-endoscopes. However, the limited availability of training data remains a significant barrier. As a result, existing research of SDE in medical imaging put extra effort into learning through view synthesis [recasens2021, shao2022self, liu2019dense, ozyoruk2021endoslam] and generative models [tian2024dd, mathew2020augmenting, shen2019context, wei2024absolute] to estimate depth for individual datasets or to enhance unsupervised training efficacy. With the success of foundation models and their outstanding performance in zero-shot SDE, a few recent works attempted to adapt foundation models trained on natural images to medical imaging [cui2024surgical, cui2024endodac]. However, due to the large domain gap between medical images and natural images, zero-shot generalization of these foundation models is not as good as expected [han2024depth].
@@ IV-A Implementation Details
We adopt the DINOv2 [oquab2024dinov2] encoder and DPT [ranftl2021vision] decoder for both the teacher and student models, as introduced by [yang2024depth]. We first train the teacher model with a ViT-L backbone [dosovitskiy2021an] on all labeled training sets until convergence. We then train the student model with a batch size of 32 until convergence under the teacher model’s guidance. We offer three student model scales based on ViT-S, ViT-B, and ViT-L architectures respectively. Following the approach in [yang2024depth], we challenge the student model with strong perturbations including color jittering and Gaussian blurring. Additionally, a semantic perception technique aligning backbone features with DINOv2 [oquab2024dinov2] is adopted. Labeled data and unlabeled data are mixed in a 1:3 ratio for each input batch. For both teacher and student network training, we initialize the encoder with pretrained weights [yang2024depth] for faster convergence and randomly initialize the decoder weights. The choice of pretrained Depth Anything weights over alternatives like Depth Anything v2 [yang2024depth2] is based on preliminary experiments. We use a learning rate of $5\times 10^{-6}$ and a learning rate factor of 0.1:1 for the encoder and decoder. The networks are trained with the Adam optimizer with a linear decay learning rate.
@@
Figure 5: Zero-shot performance of EndoOmni on SERV-CT (top two rows) and the Hamlyn Dataset (bottom two rows), compared with EndoDAC [cui2024endodac], the leading SOTA method for endoscopy, and Depth Anything [yang2024depth], the top-performing SOTA foundation model. Quantitative results are also provided for the model without our robust training loss, denoted as SSI. We show corresponding point clouds rendered from the same viewpoint. Misalignments with ground truth are highlighted in boxes on the bottom row for easy identification. EndoOmni demonstrates the most accurate geometric structure and the sharpest edges among all methods.
@@ IV-C Zero-shot Relative Depth Estimation
Compared with existing research on SDE in medical imaging, one important advantage of leveraging a foundation model trained on abundant mixed data is accurate depth estimation on unseen datasets during training. Therefore, we validate zero-shot RDE performance on two endoscopic datasets: Hamlyn and SERV-CT [edwards2020serv]. We follow existing works on MDE in medical imaging, fine-tuning the model on the SCARED dataset [allan2021stereo] at the third stage of training of our foundation model and use the same split scheme as existing works [recasens2021, cui2024surgical, cui2024endodac]. As shown in Table II and Table III, our model surpasses existing methods by a large margin. Notably, the best performing existing state-of-the-art (SOTA) method on the Hamlyn and SERV-CT Dataset, EndoDAC [cui2024endodac], focuses on efficiently adapting Depth Anything-B [yang2024depth] to endoscopic scenes through parameter-efficient fine-tuning. Compared with EndoDAC, our EndoOmni-B improves on all evaluated metrics by a large margin. These significant improvements highlight the advantages of robust learning with large, diverse data, which supports training larger models and learning more generalizable features. Our model thus gains a specialized feature understanding of medical data through our curated endoscopy datasets and advanced robust training framework. Quantitative results are shown in Fig. 5. Our EndoOmni provides the highest-quality outputs with well-defined edges.
TABLE VII: Ablation studies on baselines and losses.
Loss SERV-CT Hamlyn
Depth Anything 0.050 0.424 5.590 0.065 0.987 0.111 2.025 9.843 0.140 0.852
Depth Anything v2 0.049 0.422 5.556 0.066 0.987 0.110 1.977 9.847 0.140 0.852
SSI 0.044 0.332 5.115 0.057 0.993 0.110 1.996 9.761 0.139 0.854
TSSI 0.044 0.320 5.074 0.058 0.995 0.109 1.970 9.705 0.139 0.856
Ours: WSSI-labeled 0.042 0.295 4.853 0.054 0.995 0.108 1.936 9.569 0.137 0.857
Ours: WSSI 0.040 0.270 4.716 0.053 0.997 0.106 1.830 9.362 0.135 0.864
Figure 6: Stability of training across different loss functions.
TABLE VIII: Ablation on the size of network backbone.
@@
TABLE IX: Fine-tuned to segmentation on Kvasir dataset.
@@
TABLE X: Branch-level endoscope localization accuracy by fine-tuning to airway lumen detection.
@@ IV-E Ablation Studies
Robust Training Loss. We validate the effectiveness of our proposed robust training loss by analyzing the zero-shot RDE performance following the first two stages of training. Since the second stage employs a scale-and-shift invariant loss, we evaluate RDE by aligning the per-frame scale and shift with ground truth. We compare the performance of four different training losses: (1) SSI, the standard scale-and-shift invariant loss initially proposed by MiDaS [ranftl2020towards]; (2) TSSI, a loss proposed by [ranftl2020towards], where outliers are identified and removed by trimming the largest loss residuals per instance; (3) Ours: WSSI-labeled, which incorporates the teacher model to assess ground truth confidence, enhancing learning bias for labeled data; and (4) Ours: WSSI, which is our complete weighted scale and shift invariant loss that extends confidence analysis to both labeled and unlabeled data. We evaluate these losses on the SERV-CT and Hamlyn datasets. As shown in Table VII, our WSSI consistently outperforms other methods across all metrics on both datasets, demonstrating the robustness of WSSI in handling diverse data distributions. Interestingly, while TSSI shows minor improvements over SSI, the performance gain is marginal when compared to WSSI. This reinforces the idea that trimming loss outliers is not the most efficient approach for improving performance on medical image datasets.
Additionally, we examine the training dynamics of the different losses, as shown in Fig. 6. Both WSSI and TSSI contribute to training stability, while the standard SSI loss leads to training collapse due to label noise interference. This results in premature underfitting, with the best-performing model on the validation set producing depth maps with blurred edges, as shown in Fig. 5.
Comparison with Depth Anything. Since EndoOmni initializes encoder weights from the general foundation model Depth Anything, we compare our zero-shot RDE performance to demonstrate the improvements in endoscopy data. We also include a comparison with the latest Depth Anything v2 for reference. As shown in Table VII, our EndoOmni outperforms all baselines across the evaluated metrics, particularly on the SERV-CT dataset, which offers more accurate dense depth labels, highlighting our model’s superior performance.
Ablation on Backbone Size. To accommodate varying application requirements, we provide three student model sizes based on the ViT-S (24.8M parameters), ViT-B (97.5M parameters), and ViT-L (335.3M parameters) architectures. Their zero-shot performance results are presented in Table VIII. The results indicate that EndoOmni’s performance shows only a slight decline as the backbone parameter count decreases, making it suitable for resource-constrained environments.
@@ IV-F Fine-tuned to Polyp Segmentation
Although originally trained for the SDE task, EndoOmni is expected to transfer effectively to other downstream tasks in endoscopy. To validate this capability, we applied EndoOmni with the ViT-L backbone to the task of polyp segmentation. We followed the data split of the established benchmark [fan2020pranet] and conducted experiments on the Kvasir-SEG dataset [jha2020kvasir]. For this task, we modified EndoOmni by removing the final ReLU activation and adding batch normalization and dropout layers, without introducing any task-specific modules for segmentation. The results, shown in Table IX, demonstrate that EndoOmni achieves superior performance compared to existing methods in medical imaging segmentation, highlighting its strong semantic representation capabilities. Notably, we observed a significant improvement in Dice score from 0.883 to 0.909. Additionally, our EndoOmni adaptation for the segmentation task achieved compatible performance with Polyp-SAM-L [li2024polyp], a fine-tuned Segment Anything (SAM) [kirillov2023segment] foundation model. This observation underscores EndoOmni’s potential as a versatile foundation model for multi-task applications in endoscopy.
@@ IV-G Fine-tuned to Bronchoscopy Localization
To further evaluate EndoOmni’s transferability to other endoscopic tasks, we fine-tuned it to airway lumen detection. This problem was initially introduced by [sganga2019autonomous] and later used for branch-level endoscope localization in the airway [tian2024bronchotrack]. We hypothesize that the geometric structure semantics learned during SDE training enable EndoOmni to transfer learning to anatomical understanding effectively. We initialized the pretrained EndoOmni encoder with the ViT-L backbone and employed a 3-layer MLP to detect airway branches in each frame. The final endoscope localization accuracy was evaluated using the same test set from [tian2024bronchotrack]. As shown in Table X, the fine-tuned EndoOmni outperformed existing specialized methods for branch-level endoscope localization. Additionally, we tested initializing the encoder with the powerful DINOv2 [oquab2024dinov2] pretrained weights, and EndoOmni still outperformed with a comfortable margin. This further indicates that EndoOmni could serve as a robust foundation for other downstream tasks in endoscopy.
@@ V Discussion
Another key contribution of this work is addressing the training instability observed when using the standard SSI loss [ranftl2020towards] on medical data, resulting in depth predictions often exhibiting blurry edges (see Fig. 5, SSI column). In contrast, general-domain foundation models [yang2024depth] produce sharper edges (see Fig. 5, Depth Anything column). To tackle this, we developed a refined self-learning framework with a robust WSSI loss. Our analysis shows that clean labels converge more reliably, and pseudo-labels become more dependable when the teacher model’s predictions are consistent across augmentations. Building on this, we estimate per-pixel confidence from the teacher model’s behavior, guiding student model training with both labels and confidence. The WSSI loss dynamically adjusts learning weights based on confidence, mitigating noise while preserving challenging regions, ultimately enhancing depth estimation performance.
The TSSI loss proposed by MiDas [ranftl2020towards], by contrast, employs a trimming strategy that excludes high-loss components, which can overlook complex image regions. Our ablation studies indicate that TSSI does not consistently outperform SSI in medical imaging, as evidenced by results in Table VII. WSSI reliably outperforms both SSI and TSSI, achieving statistically significant gains on all metrics with the Hamlyn dataset.
While we curated the largest meta-dataset for endoscopic depth estimation, additional datasets could further improve generalizability. Future work will expand our dataset with more labeled and unlabeled data for training. Although this study emphasizes robustness against label noise, endoscopic data also presents unique challenges from image artifacts like occlusions, bubbles, and specular reflections. We aim to address these issues in future work by developing robust learning methods tailored to the complexities of endoscopic imaging. Additionally, while we provide a fine-tuning approach for metric depth estimation in endoscopic imaging—vital for tasks like localization and navigation [tian2024pans]—achieving zero-shot metric depth estimation across varied camera systems remains an open challenge [bhat2023zoedepth, piccinelli2024unidepth], which we leave for future exploration.

### v4

@@ I Introduction
II Related Works
SDE in Medical Endoscopy. SDE in endoscopic images aids in precise navigation, accurate lesion assessment, and improved surgical planning. Given the impracticality of using depth cameras or LiDAR in such settings, most datasets depend on correlating endoscopic views with pre-operative CT scans or generating sparse point clouds via calibrated stereo-endoscopes. However, the limited availability of training data remains a significant barrier. As a result, existing research of SDE in medical imaging put extra effort into learning through view synthesis [recasens2021, shao2022self, liu2019dense, ozyoruk2021endoslam], generative models, [tian2024dd, mathew2020augmenting, shen2019context, wei2024absolute] and Shape from Shading [puigvert2023lightdepth] to estimate depth for individual datasets by enhancing unsupervised training efficacy or to overcome illumination changes. With the success of foundation models and their outstanding performance in zero-shot SDE, a few recent works attempted to adapt foundation models trained on natural images to medical imaging [cui2024surgical, cui2024endodac]. However, due to the large domain gap between medical images and natural images, zero-shot generalization of these foundation models is not as good as expected [han2024depth].
@@ IV-A Implementation Details
We adopt the DINOv2 [oquab2024dinov2] encoder and DPT [ranftl2021vision] decoder for both the teacher and student models, as introduced by [yang2024depth]. We first train the teacher model with a ViT-L backbone [dosovitskiy2021an] on all labeled training sets until convergence. We then train the student model with a batch size of 32 until convergence under the teacher model’s guidance. We offer three student model scales based on ViT-S, ViT-B, and ViT-L architectures respectively. Following the approach in [yang2024depth], we challenge the student model with strong perturbations including color jittering and Gaussian blurring. Additionally, a semantic perception technique aligning backbone features with DINOv2 [oquab2024dinov2] is adopted. Labeled data and unlabeled data are mixed in a 1:3 ratio for each input batch. For both teacher and student network training, we initialize the encoder with pretrained weights [yang2024depth] for faster convergence and randomly initialize the decoder weights. We use a learning rate of $5\times 10^{-6}$ and a learning rate factor of 0.1:1 for the encoder and decoder. The networks are trained with the Adam optimizer with a linear decay learning rate.
@@
Figure 5: Zero-shot performance of EndoOmni on SERV-CT (top two rows) and the Hamlyn Dataset (bottom two rows), compared with EndoDAC [cui2024endodac], the leading SOTA method for endoscopy, and Depth Anything [yang2024depth], the top-performing SOTA foundation model. Quantitative results are also provided for the model without our robust training loss, denoted as SSI. We show corresponding point clouds rendered from the same viewpoint. Misalignments with ground truth are highlighted in boxes on the bottom row for easy identification.
@@ IV-C Zero-shot Relative Depth Estimation
Compared with existing research on SDE in medical imaging, one important advantage of leveraging a foundation model trained on abundant mixed data is accurate depth estimation on unseen datasets during training. Therefore, we validate zero-shot RDE performance on two endoscopic datasets: Hamlyn and SERV-CT [edwards2020serv]. We follow existing works on MDE in medical imaging, fine-tuning the model on the SCARED dataset [allan2021stereo] at the third stage of training of our foundation model and use the same split scheme as existing works [recasens2021, cui2024surgical, cui2024endodac]. As shown in Table II and Table III, our model surpasses existing methods by a large margin. Notably, the best performing existing state-of-the-art (SOTA) method on the Hamlyn and SERV-CT Dataset, EndoDAC [cui2024endodac], focuses on efficiently adapting Depth Anything-B [yang2024depth] to endoscopic scenes through parameter-efficient fine-tuning. Compared with EndoDAC, our EndoOmni-B improves on all evaluated metrics by a large margin. These significant improvements highlight the advantages of robust learning with large, diverse data, which supports training larger models and learning more generalizable features. Our model thus gains a specialized feature understanding of medical data through our curated endoscopy datasets and advanced robust training framework. Quantitative results are shown in Fig. 5. Notably, we observed that Depth Anything and its derivative, EndoDAC, perform poorly on textureless images with smooth structures, as opposed to images with complex geometry. This could be attributed to the scarcity of such scenes in the training data for general depth estimation tasks, despite their prevalence in endoscopic images. In contrast, our EndoOmni consistently delivers the highest-quality outputs.
Figure 6: Relative performance on (a) SERV-CT and (b) Hamlyn datasets, using our $L_{\text{wssi}}$ as the reference. The four losses in the white region consistently outperform the existing foundation models, Depth Anything (DA) and Depth Anything v2 (DA2), shown in the gray region. Our complete $L_{\text{wssi}}$ achieves superior performance across all metrics, surpassing both existing methods and other loss functions, owing to its robustness against label noise.
TABLE VII: Ablation on the size of network backbone.
@@
TABLE VIII: Fine-tuned to segmentation on Kvasir dataset.
@@
TABLE IX: Branch-level endoscope localization accuracy by fine-tuning to airway lumen detection.
@@ IV-E Ablation Studies
Robust Training Loss. We validate the effectiveness of our proposed robust training loss by analyzing the zero-shot RDE performance following the first two stages of training. Since the second stage employs a scale-and-shift invariant loss, we evaluate RDE by aligning the per-frame scale and shift with ground truth. We compare the performance of four different training losses: (1) SSI, the standard scale-and-shift invariant loss initially proposed by MiDaS [ranftl2020towards]; (2) TSSI, a loss proposed by [ranftl2020towards], where outliers are identified and removed by trimming the largest loss residuals per instance; (3) Ours: WSSI-labeled, which incorporates the teacher model to assess ground truth confidence, enhancing learning bias for labeled data; and (4) Ours: WSSI, which is our complete weighted scale and shift invariant loss that extends confidence analysis to both labeled and unlabeled data. We evaluate these losses on the SERV-CT and Hamlyn datasets. As shown in Fig. 6, our WSSI consistently outperforms other methods across all metrics on both datasets, demonstrating the robustness of WSSI in handling diverse data distributions. The improvement is particularly pronounced for Sq Rel, which is more sensitive to large prediction errors in the smaller and more challenging regions. While TSSI shows improvements over SSI, the performance gain is marginal when compared to WSSI. This reinforces the idea that trimming loss outliers is not the most efficient approach for improving performance on medical image datasets.
Comparison with Depth Anything. Since EndoOmni initializes encoder weights from the general foundation model Depth Anything, we compare our zero-shot RDE performance to demonstrate the improvements in endoscopy data. We also include a comparison with the latest Depth Anything v2 for reference. As shown in Fig. 6, training on a large medical dataset generally improves SDE performance across all metrics, regardless of the evaluated loss function. Notably, our EndoOmni, leveraging the WSSI loss, consistently outperforms all baselines across the evaluated metrics, particularly on the SERV-CT dataset, which offers more accurate dense depth labels, highlighting our model’s superior performance.
Ablation on Backbone Size. To accommodate varying application requirements, we provide three student model sizes based on the ViT-S (24.8M parameters), ViT-B (97.5M parameters), and ViT-L (335.3M parameters) architectures. Their zero-shot performance results are presented in Table VII. The results indicate that EndoOmni’s performance shows only a slight decline as the backbone parameter count decreases, making it suitable for resource-constrained environments.
@@ IV-F Fine-tuned to Polyp Segmentation
Although originally trained for the SDE task, EndoOmni is expected to transfer effectively to other downstream tasks in endoscopy. To validate this capability, we applied EndoOmni with the ViT-L backbone to the task of polyp segmentation. We followed the data split of the established benchmark [fan2020pranet] and conducted experiments on the Kvasir-SEG dataset [jha2020kvasir]. For this task, we modified EndoOmni by removing the final ReLU activation and adding batch normalization and dropout layers, without introducing any task-specific modules for segmentation. The results, shown in Table VIII, demonstrate that EndoOmni achieves superior performance compared to existing methods in medical imaging segmentation, highlighting its strong semantic representation capabilities. Notably, we observed a significant improvement in Dice score from 0.883 to 0.909. Additionally, our EndoOmni adaptation for the segmentation task achieved compatible performance with Polyp-SAM-L [li2024polyp], a fine-tuned Segment Anything (SAM) [kirillov2023segment] foundation model. This observation underscores EndoOmni’s potential as a versatile foundation model for multi-task applications in endoscopy.
@@ IV-G Fine-tuned to Bronchoscopy Localization
To further evaluate EndoOmni’s transferability to other endoscopic tasks, we fine-tuned it to airway lumen detection. This problem was initially introduced by [sganga2019autonomous] and later used for branch-level endoscope localization in the airway [tian2024bronchotrack]. We hypothesize that the geometric structure semantics learned during SDE training enable EndoOmni to transfer learning to anatomical understanding effectively. We initialized the pretrained EndoOmni encoder with the ViT-L backbone and employed a 3-layer MLP to detect airway branches in each frame. The final endoscope localization accuracy was evaluated using the same test set from [tian2024bronchotrack]. As shown in Table IX, the fine-tuned EndoOmni outperformed existing specialized methods for branch-level endoscope localization. Additionally, we tested initializing the encoder with the powerful DINOv2 [oquab2024dinov2] pretrained weights, and EndoOmni still outperformed with a comfortable margin. This further indicates that EndoOmni could serve as a robust foundation for other downstream tasks in endoscopy.
@@ V Discussion
Another key contribution of this work is addressing the disturbance of noisy depth labels on medical data, resulting in depth predictions often exhibiting blurry edges (see Fig. 5, SSI column). In contrast, general-domain foundation models [yang2024depth] produce sharper edges (see Fig. 5, Depth Anything column). To tackle this, we developed a refined self-learning framework with a robust WSSI loss. Our analysis shows that clean labels converge more reliably, and pseudo-labels become more dependable when the teacher model’s predictions are consistent across augmentations. Building on this, we estimate per-pixel confidence from the teacher model’s behavior, guiding student model training with both labels and confidence. The WSSI loss dynamically adjusts learning weights based on confidence, mitigating noise while preserving challenging regions, ultimately enhancing depth estimation performance.
@@ VI Conclusion
Acknowledgments
This work was supported by the Centre of AI and Robotics, Hong Kong Institute of Science and Innovation, Chinese Academy of Sciences, sponsored by InnoHK Funding, HKSAR, and partially supported by Institute of Automation, Chinese Academy of Sciences.